OpenStudy (anonymous):

Find 3 positive consecutive integers such that the square of the first, added to the last, is 8.

myininaya (myininaya):

Say these are the consecutive integers: n,n+1,n+2 Now it says the square of the first+last=8 So we have \[n^2+(n+2)=8\]

myininaya (myininaya):

See if you can solve this for n

OpenStudy (campbell_st):

3 integers x, x+1, x+2 then x^2 + (x+2)^2 = 8 expand and simplify x^2 + x+2 = 8 solve the quadratic x^2 + x - 6 = 0 find the positive solution

OpenStudy (campbell_st):

oops omit the line x^2 + (x+2)^ = 8.....

OpenStudy (anonymous):

im really confused can you draw it out?

myininaya (myininaya):

Which part is confusing you?

myininaya (myininaya):

Is it the 3 consecutive integers being n,n+1,n+2?

myininaya (myininaya):

or is it the square of the first added to the last is 8 confusing you?

myininaya (myininaya):

\[n^2+(n+2)=8\] n is the first n+2 is the last

OpenStudy (anonymous):

im going to try and find n .

myininaya (myininaya):

k

OpenStudy (anonymous):

So n= -3 and n= 2

OpenStudy (anonymous):

so i use 2

myininaya (myininaya):

First step to solve for n subtract 8 on both sides: \[n^2+n+2-8=0\] \[n^2+n-6=0\] \[(n+3)(n-2)=0\] Omg you are totally correct! go you brendahidalgo! :)

myininaya (myininaya):

So but it does say positive

myininaya (myininaya):

So n=2 is the positive answer of the two you have there

myininaya (myininaya):

So n+1=?

OpenStudy (anonymous):

yes it says three positive consecutive numbers

OpenStudy (anonymous):

so the numbers are 2, 3, and 4? because 2 squared is 4, and when you add 4 it is eight.

myininaya (myininaya):

yep 2,3,4 :)

OpenStudy (anonymous):

thank you so much :)

myininaya (myininaya):

Very good job brendahidalgo!